OpenStudy (anonymous):

Which fraction in simplest form is equivalent to zero point four? four-sixteenths eighty thirty-seconds twelve-thirds twenty-four sixtieths

OpenStudy (mitu12):

0.4 = 4/10 multiplying both numerator and denominator by 6 24 / 60 So answer (D) is correct.

OpenStudy (astrophysics):

eighty thirty - seconds? Never heard of that. But your answer is twenty - four sixteenths, that's 24/60 = 2/5 = 0.4